to determine number of beats for different set frequencies and verify result with beats= n1-n2
1
Expert's answer
2020-07-14T08:54:32-0400

The frequency of beats (also called beat frequency) can be expressed as difference in frequency of two notes that produce a beat when interfere with each other.

Assume that the frequencies we have are 300 and 301 Hz. The beat frequency is


"f_b=301-300=1\text{ Hz}."

1 Hz means one beat per second. For time t, the number of beats will be


"n=f_bt."


For another set of frequencies, the number of beats will be


"n_i=(f_{i+1}-f_i)t=f_{bi}t."